Not edited Temperature Change and Heat Capacity - Calorimetry- Three substances

water copper silver cup Case 1: A 0.780-kg silver pellet with a temperature of 85 °C is added to 0.150 kg of water in a coppeCase 2: A 0.780-kg silver pellet with a temperature of 85 °C is added to 0.150 kg of water in a copper cup of 0.61 kg. The inPLEASE ANSWER ALL PARTS OF THE QUESTION: CASE 1,2 and 3
Answer #1

cat silver only (: silver is at high temperature) heat release a siker Csiler (AT) silver released = 0.78% 234 *785-29) : 105case 3: 0.15x4186X(22:5-14) Qw = mucw (AT)w = / ow5337. 1537 Qc 3 maca (AT) = 0.54X 387 *( 225–14) Q = 1776.33 ū Heat release